OpenStudy (anonymous):

3-4(z+1)=3z-(2+z) solve for z PLZ HELP

OpenStudy (anonymous):

distribute: 3-4z-4=3-2-z combine like terms: -1-4z=1-z put all like terms on one side of the equation: -3z=2 divide -3 from both sides to get z by itself: z= -2/3